OpenStudy (anonymous):

a cylinder with a radius of 10 inches and a height of 36 inches is sfilled with water. If you drop a ball with a radius fo 7 inches into the cylinder how many cubic inches of water will still remain in the cylinder?

OpenStudy (dumbcow):

Volume of cylinder - volume of ball 3600pi - (4/3)*343pi = 9428/3 pi

jhonyy9 (jhonyy9):

so volum of cylinder = pir^2 * h =pi*10^2 *36 = 100pi *36 =3600pi and volum of ball = 4pir^3 /3 = 4pi*343/3 =1372pi /3 = 457,33pi so the reamainder water volum = 3600pi - 457,33pi = 3142,67pi cubic inches